find the first four terms of the sequence a) a[n]=n^3+1
b) a[1]=4;a[n]=3a[n-1]+1
thank you...

Answer by venugopalramana(3286)

find the first four terms of the sequence a) a[n]=n^3+1
PUT
N=1......A1=1^3+1=2
N=2......A2=2^3+1=9
N=3......A3=28
N=4......A4=65

b) a[1]=4;a[n]=3a[n-1]+1
PUT N=1
A1=4 GIVEN
N=2.....A2=3A1+1=3*4+1=13
N=3.....A3=3A2+1=3*13+1=40
N=4.....A4=3*A3+1=3*40+=121
thank you...